OpenStudy (anonymous):

Verify: [cos(-x) tan x]/[cos(TT/2-x)]=1

OpenStudy (anonymous):

\[LHS=\frac{ \cos(-x)\tan x }{ \cos(\dfrac{\pi}{2}-x) }\] Since \[\cos(-x)=\cos x\] and \[\cos(\dfrac{\pi}{2}-x)=\sin x\] \[=\frac{\cos(x)\times \dfrac{\sin x}{\cos x}}{\sin x}\] \[=\frac{\dfrac{\cos x \sin x}{\cos x}}{\sin x}\]Simplify that and you will get the answer you will be looking for.

OpenStudy (anonymous):

Wow that was very simple. Thank you so much for the help!!!